The gravitational force between two objects is 200 N when they are at a distance of 5m. Find the gravitational force between them when they are at a distance of 250 cm


1
Expert's answer
2021-01-14T10:37:46-0500
"F=G\frac{mM}{r^2}\\F'=G\frac{mM}{r'^2}=F\frac{r^2}{r'^2}
\\F'=200\frac{5^2}{2.5^2}=800\ N"